OpenStudy (anonymous):

Find the slope of the line through the pair of points. A(2, –3), P(2, 9)

OpenStudy (anonymous):

A. 0 it is not (B) C. -1/3 D. undefined

OpenStudy (anonymous):

whr r u stuck?

OpenStudy (anonymous):

i dont know the formila

OpenStudy (anonymous):

(x1, y1), (x2,y2) slope = y2-y1 / x2-x1

OpenStudy (anonymous):

so -1/3 :)

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

(9-(-3))/(2-0)=12/0=undefined

OpenStudy (anonymous):

thank you